# A hybrid electromechanical phase-field and deep learning framework for predicting fracture in dielectric nanocomposites

### v1

@@ 4 Conclusion and future work
This study introduced a hybrid computational framework that integrates electromechanical phase-field fracture modeling with deep learning–based surrogate models to predict crack propagation in boehmite-reinforced dielectric nanocomposite plates. The methodology is rooted in physics-informed simulations that account for both stiffness degradation and permittivity loss due to fracture. Two physically meaningful observables–the phase-field damage variable and the electrical potential distribution–were used as training data for ResNet-U-Net convolutional neural networks, enabling pixel-wise crack segmentation.
@@
    Experimental validation using digital image correlation (DIC) and embedded voltage sensors to benchmark model predictions against physical measurements.
@@
    Exploration of transfer learning across material systems, geometries, and loading conditions, enabling the development of universal fracture prediction models for smart materials.

### v2

@@ 4 Conclusion and future work
This study introduced a hybrid computational framework that integrates electromechanical phase-field fracture modeling with deep learning–based surrogate models to predict crack propagation in boehmite-reinforced dielectric nanocomposite plates. The methodology is rooted in high-fidelity physics-based simulations that account for both stiffness degradation and permittivity loss due to fracture. Two physically meaningful observables–the phase-field damage variable and the electric potential distribution–were used as training data for ResNet-U-Net convolutional neural networks, enabling pixel-wise crack segmentation.
@@
  • 6.
    The proposed framework achieves a significant reduction in computational cost; while a high-fidelity electromechanical phase-field simulation requires approximately 15 seconds per case on 32 CPU cores, the trained CNN completes the prediction in 0.1 s on a GPU, representing a 150-fold speed-up.
@@
    Experimental validation using digital image correlation (DIC) and embedded voltage sensors to benchmark model predictions against physical measurements and assess reliability under real-world loading conditions beyond simulation-based training.
@@
    Exploration of transfer learning across diverse geometry classes (e.g., notches and varying aspect ratios) and loading conditions, enabling the development of universal fracture prediction models for smart materials that generalize beyond the specific patterns used in this study.
  • 6.
    Invesitigation of the impact of alternative electrical boundary conditions and crack-face properties–such as perfectly insulating or partially conducting cracks–as discussed in recent multiferroic phase-field studies Tan et al. [2023]. Additionally, for more complex multi-crack topologies, we intend to implement granular spatial-residual error maps to further investigate the local architectural limitations of the surrogate model.